### 16 Mar 2026 v1

# Ringdown waves from hairy black holes

Abstract
We derive general formulas for quasi-normal mode (QNM) frequencies of hairy black holes by exploiting the QNM–geodesic correspondence. The black hole hair is treated as an anisotropic fluid perturbatively added to the vacuum black holes (Schwarzschild and Kerr black holes). Under this setting, independent of energy conditions, our formulas offer a systematic method to compute quasi-normal mode frequencies for a broad class of hairy black holes.
1 Introduction
Recent cosmological and astrophysical observations have accumulated compelling evidence for the existence of black holes, making them prime candidates for testing general relativity. On the theoretical side, a variety of non-vacuum or hairy black hole models have been proposed. These models may be interpreted as effective descriptions for black holes surrounded by dark sector fields or effects of modified gravity theories near a black hole. It is highly relevant to investigate the basic characteristics of hairy black holes and to develop techniques for identifying them, for instance through gravitational wave measurements. When a black hole is perturbed, the remnant emits gravitational radiation known as the ringdown phase. This signal is primarily characterized by a collection of damped oscillations whose complex frequencies correspond to the quasinormal modes (QNMs). These modes encode essential information about the black hole’s physical properties. If QNMs from a black hole endowed with “dark hair” could be detected, they would allow us to determine not only the mass and angular momentum of the source but also the state parameters of the surrounding dark-sector fields or the coupling constants that define the underlying modified gravity theory.
However, apart from a few particular cases (e.g., analysis of QNMs from Bardeen black hole by Konoplya et. al. [1]), so far, a systematic study of QNMs applicable to a broad class of hairy black holes still remains unexplored. This is in part because the metrics describing hairy black hole models are generally complicated, making it challenging to perform QNM calculations. There is a strong need to develop general analytic formulas that can be uniformly used to analyze QNMs of various types of hairy black holes, independently of the specific details of each hairy black hole model (e.g., independent of the dark sector equation of state).
To address this problem, in this paper, we consider hairy black hole models constructed by perturbatively adding anisotropic fluids (mimicking dark sector fields) to vacuum black hole metrics (i.e., the Schwarzschild and Kerr metrics). To systematically understand how QNMs of these constructed hairy black holes are modified by their non-trivial hair, compared to their corresponding vacuum counterparts, we exploit the relationship between the unstable circular orbit of a photon (UCOP) and second-order WKB results for QNMs. This relationship is well-established in the eikonal limit, where the perturbation’s angular momentum is large, but remains valid even at lower angular momentum quantum number $\ell\gtrsim 4$. Along with $\ell$, the QNMs are indexed by the quantum numbers $\mathbf{m}$, the azimuthal shape and $n$, the overtone index. In the non-rotating vacuum case, the sign difference of $\mathbf{m}$ is degenerate. This correspondence has been justified at least for static, spherically symmetric, asymptotically flat spacetimes [2]. For further justification using the Penrose limit, see Refs. [4, 3, 5]. This approach offers a significant advantage because the null geodesic equations are considerably easier to analyze than the wave equations. According to the analysis done by Cardoso et al. [2], the QNM frequency $\omega$ can be decomposed into its oscillation frequency and a decay rate. In the eikonal limit, this is determined by the orbital frequency $\Omega$ and the Lyapunov exponent $\lambda$ of UCOP,
$\omega_{\rm QNM}=\Omega\ell-i(n+1/2)\lambda\,.$ (1.1)
In this paper, we derive such a general formula for the QNMs of static spherically symmetric hairy black holes without resorting to any specific models of black hole hair. We then express $\Omega$ and $\lambda$ explicitly in terms of the state parameter of the dark sector field. In this context, it is worth noting the no-short hair theorem, according to which, if a static black hole supports hair in the form of an anisotropic fluid satisfying certain energy conditions, that hair must extend beyond the radius of the UCOP, namely the photon sphere [6].
We also perform a similar analysis for stationary rotating hairy black holes. Rotation breaks the $\mathbf{m}$ degeneracy and therefore, the correspondence between QNMs and UCOP does not hold in general. However it does hold for orbital rays trapped in the equatorial plane of Kerr-Newman black holes, expressed as the $\ell=|\mathbf{m}|$ modes [14, 15, 16, 21].
Recently, there have appeared a series of papers by Igata [7, 8, 9], which share some overlap with the present work. While those papers put more emphasis on the geometric aspects of UCOP and QNMs, our present paper focuses more on the relationship between QNMs and the equation of state for matter fields surrounding black holes.
The paper is organized as follows: In the next section, we derive the UCOP for a general static spherically symmetric black hole geometry. We also discuss the relation between UCOP and QNM. In section 3, as our examples of hairy static black holes, we examine the cases of Bardeen, Hayward, and Kiselev metrics. In section 4, we derive general formulas for stationary rotating hairy black holes. In section 5, we examine the rotating version of the models previously introduced in section 3. In the appendix, we describe some useful geometric formulas for our analyses. This paper uses the geometrized units convention ($G=1$, $c=1$).
2 Static hairy black holes and QNMs
In this section, we evaluate the components ($\Omega$, $\lambda$) of the QNMs (1.1) for static spherical hairy black holes from the orbital frequency and Lyapunov exponent for the UCOP.
2.1 Static hairy black holes and null geodesics
We start with the following general static spherically symmetric metric
$ds^{2}=-f(r)dt^{2}+h(r)dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2})\,.$ (2.1)
[…]
${\cal L}:=\dfrac{1}{2}\left(-f\dot{t}^{2}+h\dot{r}^{2}+r^{2}\dot{\varphi}^{2}\right)=\dfrac{\epsilon}{2}\,$, (2.2)
where we have already set $\theta=\pi/2$ and $\epsilon=-1$ for timelike geodesics and $\epsilon=0$ for null geodesics. From the isometries along $t$ and $\varphi$ we immediately find the two conserved quantities
$E:=-\dfrac{\partial{\cal L}}{\partial\dot{t}}=f\dot{t}\,,\quad L:=\dfrac{\partial{\cal L}}{\partial\dot{\varphi}}=r^{2}\dot{\varphi}\,.$ (2.3)
Plugging these two into the Lagrangian (3.6), we obtain
$\dfrac{1}{2}\dot{r}^{2}+V(r)=0\,,\quad V(r):=\dfrac{1}{2h}\left(-\epsilon+\dfrac{L^{2}}{r^{2}}-\dfrac{E^{2}}{f}\right)\,.$ (2.4)
The conditions for geodesic curves to be spatially closed are given by
$V=0\,,\quad V^{\prime}=0\,$, (2.5)
on the orbit, where the prime denotes the derivative by $r$. From the above two conditions, we obtain, respectively,
$\dfrac{L^{2}}{r^{2}}=\dfrac{E^{2}}{f}+\epsilon\,,\quad\dfrac{2L^{2}}{r^{3}}=\dfrac{f^{\prime}}{f^{2}}E^{2}\,.$ (2.6)
Combining these two, we obtain
$\epsilon\circeq\dfrac{E^{2}}{2f^{2}}(rf^{\prime}-2f)\,$, (2.7)
[…]
$V^{\prime\prime}\circeq\dfrac{1}{2h}\left[6\dfrac{L^{2}}{r^{4}}+\left(\dfrac{f^{\prime}}{f^{2}}\right)^{\prime}E^{2}\right]\,.$ (2.8)
Let us focus on null geodesic curves describing UCOP. Combining Eqs. (2.7) and (2.8) with $\epsilon=0$, we have
$V^{\prime\prime}\circeq\dfrac{L^{2}}{2r^{2}h}\left(\dfrac{f^{\prime\prime}}{r}-\dfrac{2}{r^{2}}\right)\,.$ (2.9)
Then, following the reference [2], we obtain the Lyapuonv exponent $\lambda$ and the orbital frequency $\Omega$ as
$\lambda^{2}:=-\dfrac{V^{\prime\prime}}{\dot{t}^{2}}$ $\circeq$ $\dfrac{f}{2h}\left(\dfrac{2}{r^{2}}-\dfrac{f^{\prime\prime}}{f}\right)\,$, (2.10)
$\Omega:=\dfrac{\dot{\varphi}}{\dot{t}}$ $\circeq$ $\dfrac{\sqrt{f}}{r}\,.$ (2.11)
For the Schwarzschild metric case, i.e., $f=h^{-1}=1-2M/r$, we obtain the well-known result
$\lambda_{0}=\dfrac{1}{3\sqrt{3}M}\,,\quad\Omega_{0}=\dfrac{1}{3\sqrt{3}M}\,.$ (2.12)
2.2 Anisotropic fluid surrounding black holes and QNM frequency
Let us consider the case in which our metric solves the Einstein equations. We consider fluids with anisotropic pressure so that the stress-energy tensor $T^{\mu}{}_{\nu}$ is given by the components,
$T^{t}{}_{t}=-\rho\,,\quad T^{r}{}_{r}=P_{r}\,,\quad T^{\theta}{}_{\theta}=T^{\varphi}{}_{\varphi}=P_{\theta}\,$, (2.13)
and the rest of the components are vanishing. Combining the components of the Einstein tensor, $G_{00}=8\pi\rho$, $G^{r}{}_{r}=8\pi P_{r}$, and $G^{\theta}{}_{\theta}=8\pi P_{\theta}$, we obtain
$\lambda^{2}\circeq\dfrac{f}{r^{2}}-\dfrac{f}{r^{2}}\left[2\pi r^{2}(\rho-3P_{r}+4P_{\theta})+2\pi r^{2}h(\rho+P_{r})+16\pi^{2}r^{4}hP_{r}(\rho+P_{r})\right]\,.$ (2.14)
Now, let us suppose that the equations of state for our fluid are given by two parameters
$P_{r}=w_{r}\rho\,,\quad P_{\theta}=w_{\theta}\rho\,$, (2.15)
and furthermore that the anisotropic fluid can be treated as linear perturbations from the vacuum Schwarzschild background, so that
$|\rho|\ll 1\,.$ (2.16)
Then, denoting the UCOP radius of the hairy black hole by $r_{\star}$, and approximating $h^{-1}(r)\approx f_{0}(r)=1-2M/r$, justified by the fact that it appears multiplied by the perturbative fluid density $\rho$, we obtain the leading-order expression for $\lambda$ (2.14) at the UCOP location as follows
$\lambda_{\star}$ $\simeq$ $\Omega_{\star}-\Omega_{\star}\pi r_{\star}^{2}\rho\left[1-3\,w_{r}+4\,w_{\theta}+\dfrac{(1+w_{r})}{f_{0}}\right]\bigg|_{r=r_{\star}}\,$, (2.17)
where $\Omega_{\star}=\dfrac{\sqrt{f(r_{\star})}}{r_{\star}}$ represents the orbital frequency at the UCOP.
To obtain fully linearized expressions for the QNM coefficients around the Schwarzschild background, we further need to expand
$r_{\star}\simeq r_{0}+\delta r=3M+\delta r\,$, (2.18)
$f(r)\simeq f_{0}(r)+\delta f(r)\,$, (2.19)
$\lambda_{\star}\simeq\lambda_{0}+\delta\lambda=(3\sqrt{3}M)^{-1}+\delta\lambda\,$, (2.20)
$\Omega_{\star}\simeq\Omega_{0}+\delta\Omega=(3\sqrt{3}M)^{-1}+\delta\Omega\,.$ (2.21)
where $r_{0}=3M$ and $f_{0}(r_{0})=1/3$ denote the Schwarzschild UCOP radius and the corresponding metric function value, respectively. The terms $\delta r$ and $\delta f$ represent small linear deviations from these Schwarzschild quantities.
We start by substituting values in the UCOP expression $g(r)=rf^{\prime}(r)-2f(r)$, derived from (2.7), which vanishes for the circular photon orbit. By evaluating it in the UCOP radius for the hairy black hole, we obtain
$g(r_{\star})\simeq((r_{0}+\delta r)(f_{0}^{\prime}(r)+\delta f^{\prime}(r))-2(f_{0}(r_{0}+\delta r)+\delta f(r_{0}+\delta r)))|_{r=r_{0}+\delta r}\,.$ (2.22)
By using the result for the Schwarzschild photon sphere $g(r_{0})=r_{0}f_{0}^{\prime}(r_{0})-2f_{0}(r_{0})=0$, and by keeping perturbation terms up to the first order, we can solve this expression for $\delta r$ as
$\delta r\simeq\frac{r_{0}\,\delta f^{\prime}(r_{0})-2\,\delta f(r_{0})}{f_{0}^{\prime}(r_{0})-r_{0}f_{0}^{\prime\prime}(r_{0})}\,=\,\frac{3M}{2}(r_{0}\,\delta f^{\prime}(r_{0})-2\,\delta f(r_{0}))\,.$ (2.23)
In a similar way, we proceed to linearize the orbital frequency (2.11) evaluated at the UCOP
$\Omega_{\star}^{2}\simeq\frac{1}{3r_{0}^{2}}(1+3\,\delta f(r_{0}))\,$, (2.24)
which leads to
$\Omega_{\star}\simeq\Omega_{0}(1+\frac{3}{2}\delta f(r_{0}))\,.$ (2.25)
Finally, by substituting $\Omega_{\star}$ into the Lyapunov exponent (2.17), we obtain
$\lambda_{\star}\simeq\frac{1}{\sqrt{3}\,r_{0}}(1+\frac{3}{2}\delta f)\,\bigg(1-\pi r_{\star}^{2}\rho(1-3w_{r}+4w_{\theta}+\frac{(1+w_{r})}{f_{0}})\bigg)\bigg|_{r=r_{0}}\,$, (2.26)
and by expanding up to the first order for the perturbations, we obtain
$\lambda_{\star}\simeq\lambda_{0}(1-4\pi r_{0}^{2}\rho(1+w_{\theta})+\frac{3}{2}\delta f)|_{r=r_{0}}\,.$ (2.27)
2.3 Modifications of QNMs and Energy Conditions
In this section, we quantify the deviation of the QNMs for hairy black holes from their Schwarzschild counterparts, and relate them to the fluid state parameters $w_{r}$ and $w_{\theta}$. We can observe how it affects the QNM frequency as
$\frac{\delta\Omega}{\Omega_{0}}=\frac{3}{2}\,\delta f(r_{0}),\qquad\frac{\delta\lambda}{\lambda_{0}}=\frac{3}{2}\,\delta f(r_{0})-4\pi r_{0}^{2}\rho\big[1+w_{\theta}\big]|_{r=r_{0}}.$ (2.28)
Both shifts are controlled by the metric correction $\delta f(r_{0})$ in the same way, but the Lyapunov exponent also presents an explicit contribution from the tangential pressure $P_{\theta}$.
Although $\delta f$ appears geometric, it is not independent of the matter. To illustrate this, we re-parametrize our general metric as
$f(r)=1-\frac{2m(r)}{r},\quad h(r)=\left(1-\frac{2m(r)}{r}\right)^{-1},$ (2.29)
where $m(r)\simeq M+\delta m(r)$. At first order expansion we obtain an explicit expression for the geometric term
$\delta f(r)=-\frac{2\,\delta m(r)}{r}.$ (2.30)
By using the Einstein equations, we can relate the metric function $m(r)$ to the matter field as
$m^{\prime}(r)=\delta m^{\prime}(r)=4\pi r^{2}\rho(r)$ (2.31)
so that $\delta f$ is ultimately determined by $\rho$. Assuming $m(\infty)=M$, we can rewrite the QNM shifts directly in terms of the fluid parameters as follows
$\frac{\delta\Omega}{\Omega_{0}}\simeq\frac{12\pi}{r}\int_{r}^{\infty}\rho(s)s^{2}\,ds\,|_{r=r_{0}},$ (2.32)
$\frac{\delta\lambda}{\lambda_{0}}\simeq\frac{12\pi}{r}\int_{r}^{\infty}\rho(s)s^{2}\,ds-4\pi r_{0}^{2}\rho(1+w_{\theta})|_{r=r_{0}}\,.$ (2.33)
As shown in Appendix C, the conservation of an anisotropic fluid, expressed by the relation $\nabla_{\mu}T^{\mu}_{\nu}=0$, leads to a generalized Tolman–Oppenheimer–Volkoff (TOV) equation. This relation connects the radial and tangential pressures as
$P_{\theta}(r)=P_{r}(r)+\frac{r}{2}P^{\prime}_{r}(r)+\frac{(P_{r}+\rho)}{2}\left(\frac{4\pi r^{3}P_{r}(r)+m(r)}{r-2m(r)}\right)\,$, (2.34)
which shows that $P_{\theta}$ is constrained once $\rho$ and $w_{r}$ are specified.
Energy conditions require normal and healthy properties of the matter field from a viewpoint of certain causality or stability. It is worthwhile to interpret the relationship between the properties of matter fields surrounding a black hole and QNMs in terms of energy conditions. The most used conditions in general relativity are the Null, Weak, Strong, and Dominant energy conditions, which are locally defined and often abbreviated as: NEC, WEC, SEC, and DEC, respectively. WEC and SEC imply NEC, while DEC implies WEC. For their definitions and usages, see, e.g, a review [20] and references therein. Satisfying these conditions is necessary for interpreting the source as regular matter. For an anisotropic effective fluid with energy density $\rho$ and principal pressures $(P_{r},P_{\theta},P_{\theta})$, the energy conditions then read
NEC: $\quad\rho+P_{r}\geq 0,\ \ \rho+P_{\theta}\geq 0\ ,$ (2.35)
WEC: $\quad\rho+P_{r}\geq 0,\ \ \rho+P_{\theta}\geq 0\ \&\ \ \rho\geq 0\ \ \Longleftrightarrow\ \rho\geq 0,\ w_{r}\geq-1,\ w_{\theta}\geq-1,$ (2.36)
SEC: $\quad\rho+P_{r}\geq 0,\ \ \rho+P_{\theta}\geq 0\ \ \&\ \ \rho+P_{r}+2P_{\theta}\geq 0\ ,$ (2.37)
DEC: $\quad|P_{r}|\leq\rho,\ \,|P_{\theta}|\leq\rho\,\Longleftrightarrow\,\rho\geq 0\ ,\ \ |w_{r}|\leq 1,\ |w_{\theta}|\leq 1\,.$ (2.38)
In our barotropic model of an anisotropic fluid, with $i$ as an indicator of radial or tangential, the parameter $w_{i}=0$ correspond to dust, $w_{i}=1/3$ to radiation, $w_{i}<-1/3$ to dark-energy and $w_{i}<-1$ to phantom energy. A particularly important case is $w_{r}=-1$ (cosmological-constant-like). Not only does it satisfy most energy conditions, but it is also the only scenario that ensures the continuity of the energy density across the black hole event horizon [10] with non-zero energy density. In this scenario $\rho+P_{r}=0$ (at least on the horizon), and equation (2.34) reduces to
$w_{\theta}=-1-\frac{r}{2}\frac{\rho^{\prime}(r)}{\rho(r)}=-\frac{r}{2}\frac{m^{\prime\prime}(r)}{m^{\prime}(r)}.$ (2.39)
By considering the expression (2.39), we can derive an expression for the radial pressure in terms of the second derivative of the mass function $P_{r}(r)=-\frac{m^{\prime\prime}}{8\pi r}$. When considering $\rho(r)>0$, for regular matter, we find $\frac{\delta\Omega}{\Omega_{0}}>0$ from Eq. (2.32), which means the QNM of the hairy black hole oscillates faster than in the vacuum case. In particular, it is also interesting to analyze
$\frac{\delta\lambda}{\lambda_{0}}-\frac{\delta\Omega}{\Omega_{0}}\simeq-4\pi r_{0}^{2}\rho(1+w_{\theta})|_{r=r_{0}}\,.$ (2.40)
Since $\rho\,(1+w_{\theta})\geq 0$ is required by NEC (and therefore by WEC, SEC, and DEC), the difference (2.40) is negative when treating regular matter ($\rho>0$) satisfying the energy conditions. This result indicates that, if we observe a positive value of $\delta\lambda/\lambda_{0}-\delta\Omega/\Omega_{0}$, the frequency shift cannot be explained by the distribution of a regular matter field surrounding the black hole, given that our setting is relevant.
It is also interesting to examine the displacement of the UCOP radius position
$\delta r\simeq-12\,\pi\,M\left(r_{0}^{2}\rho(r)+3\int_{r}^{\infty}\rho(s)s^{2}\,ds\right)\,$, (2.41)
which, for regular matter with positive energy density, is always negative. The UCOP radius position comes closer to the center when considering the presence of regular matter hair.
3 Examples of Static Spherical Hairly Black Holes
In this section, we apply the formulas derived in the previous section to three specific black hole models: Bardeen, Hayward, and Kiselev black holes. These are asymptotically flat, static, spherically symmetric black holes that resemble the Schwarzschild spacetime at large distances, but have no singularity at the center. They are non-vacuum solutions, and hence can be interpreted as “hairy” black hole deviating from the Schwarzschild vacuum black hole.
In this section, we apply the formulas derived in the previous section to three specific static, spherically symmetric black hole models: the Bardeen, Hayward, and Kiselev spacetimes. The Bardeen and Hayward solutions are regular black holes that approach the Schwarzschild geometry at large distances. The Kiselev solution, in contrast, represents a non-vacuum black hole surrounded by an anisotropic fluid. Unlike the Bardeen and Hayward models, it typically features a central curvature singularity, and its asymptotic behaviour is determined by the chosen value of the equation-of-state parameter.
3.1 Bardeen black hole
Let us consider the Bardeen black hole, whose metric is given by (2.3) with the following components
$f(r)=1/h(r)=1-\dfrac{2Mr^{2}}{(r^{2}+q^{2})^{3/2}}\,$, (3.1)
where $q$ is a constant parameter. Bardeen metric is interpreted as the solution of the Einstein equations with a certain type of nonlinear electromagnetic source (see e.g., [11, 12]), which allows the fluid expression as
$\rho=-P_{r}=\dfrac{6Mq^{2}}{8\pi(r^{2}+q^{2})^{5/2}}\,,\quad P_{\theta}=\dfrac{Mq^{2}(9r^{2}-6q^{2})}{8\pi(r^{2}+q^{2})^{7/2}}\,.$ (3.2)
We can read off the state parameters as
$w_{r}=-1\,,\quad w_{\theta}=\dfrac{3r^{2}-2q^{2}}{2(q^{2}+r^{2})}\,.$ (3.3)
If we assume that the extra-parameter $q^{2}$ is sufficiently small, i.e., $|q/M|\ll 1$, we can treat the Bardeen solution as a perturbation from the Schwarzschild case and use the linearized expressions (2.27) and (2.25). First, we estimate $\delta f(r)$ by expanding $f(r)$ up to first order
$f(r)\simeq 1-\frac{2M}{r}\left(1-\frac{3q^{2}}{2r^{2}}\right)\,\Longrightarrow\,\delta f(r)\simeq\frac{3Mq^{2}}{r^{3}}\,.$ (3.4)
Then we obtain
$\Omega_{\star}\simeq\frac{1}{3\sqrt{3}M}\left(1+\frac{3}{2}\left(\frac{3Mq^{2}}{27M^{3}}\right)\right)\,=\,\frac{1}{3\sqrt{3}M}\left(1+\frac{q^{2}}{6M^{2}}\right)\,.$ (3.5)
By keeping terms up to $q^{2}$, we approximately obtain $4\pi r_{0}^{2}(\rho(r_{0})+P_{\theta}(r_{0}))\simeq\frac{5q^{2}}{18M^{2}}$, then
$\lambda_{\star}\simeq\frac{1}{3\sqrt{3}M}\left(1-4\pi(9M^{2})(\rho+P_{\theta})+\frac{3}{2}\left(\frac{3Mq^{2}}{27M^{3}}\right)\right)=\frac{1}{3\sqrt{3}M}\left(1-\frac{q^{2}}{9M^{2}}\right).$ (3.6)
Additionally, we can compute the shift in the photon sphere radius $\delta r$ by using equation (2.23) as
$\delta r\simeq\frac{3M}{2}\left(-\frac{q^{2}}{3M^{2}}-\frac{6q^{2}}{27M^{2}}\right)=-\frac{5q^{2}}{6M}\,$, (3.7)
which agrees with the WKB analysis of the Bardeen black hole case up to the first order [1].
In particular for this case, $\rho+P_{\theta}=\frac{15Mq^{2}r^{2}}{8\pi(q^{2}+r^{2})^{7/2}}$, then the energy conditions NEC and WEC are satisfied. Within our perturbative approximation, SEC and DEC reduce to
SEC: $\quad|q|\leq\sqrt{\frac{3}{2}}r\,\overset{\circ}{\Longrightarrow}\,\frac{|q|}{M}\leq\frac{3\sqrt{6}}{5},$ (3.8)
DEC: $\quad|q|\geq\frac{r}{2}\,\overset{\circ}{\Longrightarrow}\,\frac{|q|}{M}\geq\frac{3}{5}(\sqrt{14}-2),$ (3.9)
[…]
(a) Tangential pressure parameter $w_{\theta}(r)$ as a function of radius for various values of the charge-like parameter $q$.
(b) Relative shifts of the QNM frequancy $\delta\Omega/\Omega_{0}$ and damping rate $\delta\lambda/\lambda_{0}$ as functions of $q/M$.
Figure 1: Tangential pressure parameter $w_{\theta}(r)$ and QNM shifts for the Bardeen black hole.
In figure 1(a), we observe that the tangential fluid parameter $w_{\theta}$ becomes smaller if we increase the value of the hairy parameter $q$. At the asymptotic region, the value of the tangential parameter tend to $1.5$ for any finite value of $q$. In figure 1(b) we can observe that the $\delta\Omega/\Omega_{0}$ is positive and $\delta\lambda/\lambda_{0}$ negative as expressed in (3.5) and (3.6). Where $\delta\Omega/\Omega_{0}$ would be positive and $\delta\lambda/\lambda_{0}$ would be negative for any value of $q$. Furthermore, equation (3.2) shows that the mass density $\rho$ remains positive for all values of $q$.
3.2 Hayward black hole
The metric function of the Hayward black hole is given by
$f(r)=1/h(r)=1-\dfrac{2Mr^{2}}{r^{3}+q^{3}}\,$, (3.10)
where $q$ is a constant.
Through the Einstein equations, we can write down the energy density and pressures in terms of $M$ and $q$ as
$8\pi\rho=\dfrac{6q^{3}M}{(r^{3}+q^{3})^{2}}\,,\quad P_{r}=-\rho\,,\quad 8\pi P_{\theta}=-\dfrac{6q^{3}M(q^{3}-2r^{3})}{(r^{3}+q^{3})^{3}}\,.$ (3.11)
We can read off the state parameters as
$w_{r}=-1\,,\quad w_{\theta}=-\dfrac{q^{3}-2r^{3}}{q^{3}+r^{3}}\,.$ (3.12)
In the scenario where the extra-parameter $q^{3}$ is sufficiently small, i.e., $|q/M|\ll 1$, around the UCOP $r_{\star}$, let’s estimate $\delta f(r)$ by expanding $f(r)$ up to first order
$f(r)\simeq 1-\frac{2M}{r}\left(1-\frac{q^{3}}{r^{3}}\right)\,\Longrightarrow\,\delta f(r)\simeq\frac{2Mq^{3}}{r^{4}}.$ (3.13)
We can calculate the quasinormal mode components. By keeping terms up to $q^{3}$, we obtain $4\pi r_{0}^{2}(\rho+P_{\theta})\simeq\frac{q^{3}}{9M^{3}}$ and
$\Omega_{\star}\simeq\frac{1}{3\sqrt{3}M}\left(1+\frac{q^{3}}{27M^{3}}\right)\,,\quad\lambda_{\star}\simeq\frac{1}{3\sqrt{3}M}\left(1-\frac{2q^{3}}{27M^{3}}\right).$ (3.14)
Additionally, we can compute the shift in the photon sphere radius $\delta r$ by using equation (2.23) as
$\delta r\simeq-\frac{2q^{3}}{9M^{2}}$ (3.15)
(a) Tangential pressure parameter $w_{\theta}(r)$ as a function of radius for various values of the charge-like parameter $q$.
(b) Relative shifts of the QNM frequency $\delta\Omega/\Omega_{0}$ and damping rate $\delta\lambda/\lambda_{0}$ as functions of $q/M$.
Figure 2: Tangential pressure parameter $w_{\theta}(r)$ and frequency shifts for the Hayward black hole.
In particular for this case, $\rho+P_{\theta}=\frac{6q^{3}M}{8\pi(q^{3}+r^{3})^{3}}$, then the energy conditions NEC and WEC are accomplished when $q>0$. SEC and DEC reduce to
SEC: $\quad q\leq-r\quad\text{or}\quad 0\leq q\leq\sqrt[3]{2}\,r\,\overset{\circ}{\Longrightarrow}\,\begin{aligned} &0\leq\frac{q}{M}\leq\bigg({\frac{27}{4}+\frac{\sqrt{189}}{2}}\bigg)^{1/3}+\bigg({\frac{27}{4}-\frac{\sqrt{189}}{2}}\bigg)^{1/3}\\ &\text{or}\quad\frac{q}{M}>3\end{aligned}$ (3.16)
DEC: $\quad q>\frac{r}{\sqrt[3]{2}}\,\overset{\circ}{\Longrightarrow}\,\frac{q}{M}>\bigg({\frac{27+\sqrt{837}}{4}}\bigg)^{1/3}+\bigg({\frac{27-\sqrt{837}}{4}}\bigg)^{1/3}.$ (3.17)
As well as the previous case, NEC, WEC, and SEC are compatible with $|q|/M\ll 1$, and the Dominant condition is not compatible with a small $|q|/M\ll 1$.
Since positive values of $q$ are the only ones who secure that the mass density $\rho$ remains positive as seen in equation (3.11), let us focus on the cases with $q>0$. In figure 2(a), we observe that the tangential fluid parameter $w_{\theta}$ becomes smaller at the UCOP if we increase the value of the hairy parameter $q$. At the asymptotic region, the value of the tangential parameter tends to $2$ for any finite value of $q$. In figure 2(b) we can observe that the $\delta\Omega/\Omega_{0}$ is positive and $\delta\lambda/\lambda_{0}$ negative as expressed in (3.14).
3.3 Kiselev black hole
The Kiselev solution [13] describes static spherically symmetric black holes with quintessential matter distribution. The metric function is given by
$f(r)=1/h(r)=1-\dfrac{2M}{r}-\dfrac{k}{r^{1+3\,w_{q}}}\,$, (3.18)
where $k$ controls the strength of the surrounding matter field and $w_{q}$ is an effective state parameter. In the original quintessence interpretation, one usually considers the range $1<w_{q}<-1/3$, which is associated with accelerated expansion.
From the Einstein equations, the associated energy density and pressures are obtained as
$8\pi\rho=-\dfrac{3k\,w_{q}}{r^{3(1+w_{q})}}\,,\quad P_{r}=-\rho\,,\quad 8\pi P_{\theta}=-\dfrac{3\,k\,w_{q}(1+3\,w_{q})}{2r^{3(1+w_{q})}}\,.$ (3.19)
Accordingly, the state parameters are
$w_{r}=-1\,,\qquad w_{\theta}=\dfrac{1+3\,w_{q}}{2}\,.$ (3.20)
Since the source is anisotropic, we define the averaged pressure
$P_{q}:=\frac{P_{r}+2P_{\theta}}{3}=w_{q}\rho\,$, (3.21)
which shows that $w_{q}$ can be interpreted as the effective state parameter of the anisotropic source.
If we treat $|k|$ as a small parameter, we can treat the last term of the metric function $f(r)$ as a small deviation from the Schwarzschild metric function
$\delta f(r)=-\frac{k}{r^{1+3w_{q}}}.$ (3.22)
The frequency components are
$\Omega_{\star}\simeq\frac{1}{3\sqrt{3}M}\left(1-\frac{3k}{2(3M)^{1+3w_{q}}}\right)\,,\quad\lambda_{\star}\simeq\frac{1}{3\sqrt{3}M}\left(1+\frac{3w_{q}(1+w_{q})-2}{4(3^{3w_{q}}M^{1+3w_{q}})}k\right).$ (3.23)
Figure 3: QNM shifts for different values of $w_{q}$. Continuous lines refer to $\delta\Omega/\Omega_{0}$ and dashed lines refer to $\delta\lambda/\lambda_{0}$ for different values of $k$.
Different $w_{q}$ values tell how fast the “hair” $\delta f(r)$ decays (or grows). In its original construction, the source is “quintessence-like”, giving a negative tangential pressure. For $w_{q}=-1$, the perturbation term behaves as the cosmological constant, and both pressures are equal. For values larger than 1/3-1/3, the hair $\delta f$ decays rapidly as $r$ grows, giving us a metric more similar to the Kerr case. In particular, for the value $w_{q}=1/3$, if we consider $k=-Q^{2}$ as charge, the metric reduces to the Reissner-Nordström model. In this last case, tangential and radial pressure are equal in magnitude but opposite in sign. For the $k<0$ scenario, at figures 3(a) and 3(c), we can observe that in certain ranges of $w_{q}$, the behaviour of the QNM components resembles the Bardeen (figure 1(b)) and Hayward (figure 2(b)) scenarios: $\delta\Omega/\Omega_{0}$ becomes positive and $\delta\lambda/\lambda_{0}$ turns negative. In contrast, the “quintessence” regime, $w_{q}$ from 1-1 to 1/3-1/3, has a different behaviour as seen in figure 3(b). We observe that the sign depends on $k$, but regardless of the sign of $k$, both shifts have the same sign. In this model, $w_{\theta}$ depends only linearly on $w_{q}$.
Fig. 5 shows trajectories describing particular cases of the QNMs in the complex plane normalized by their modes. In figures 4(a), 4(b) and 4(c), the color changes mark the different values of $w_{q}$. We can observe completely different behaviors. In panel 4(d), we can observe a comparison for different values of $k$.
Additionally, we can compute the shift in the photon sphere radius $\delta r$ by using equation (2.23) as
$\delta r\simeq\frac{3k(1+w_{q})}{2(3M)^{3w_{q}}}\,.$ (3.24)
(a) Trajectory in the complex plane of a quasi-normal mode (QNM) frequency for k=0.04k=-0.04 evaluated across [2,1][-2,-1] values of the frequency state parameter $w_{q}$.
(b) Trajectory in the complex plane of a quasi-normal mode (QNM) frequency for k=0.04k=-0.04 evaluated across [1,1][-1,1] values of the frequency state parameter $w_{q}$.
(c) Trajectory in the complex plane of a quasi-normal mode (QNM) frequency for k=0.04k=-0.04 evaluated across $[1,2]$ values of the frequency state parameter $w_{q}$.
(d) Trajectory of QNM frequencies in the complex plane under varying anisotropic fluid parameter K=0.10,0.05, 0.05, 0.10K={-0.10,\,-0.05,\,0.05,\,0.10}, illustrating how the real and imaginary components of the QNM frequency shift relative to the reference point $\{\Omega_{0},\Lambda_{0}\}$ at $k=0$.
Figure 4: Representative QNM and geometric shifts for the anisotropic fluid halo Kiselev model.
(a) Damping oscillation waveforms for the Bardeen, Hayward and Schwarschild cases with $q=0.4$.
(b) Damping oscillation waveforms for the Kiselev with $k=0.04$ for different values of $w_{q}$.
Figure 5: Time evolution of the ringdown waveform, $\Psi(t)=\Re\!\left[e^{-i\,\omega_{\rm QNM}t}\right]$. We compare the amplitudes within the different models for the quasinormal mode $n=0$ and $\ell=4$ (Eikonal limit). Although $\ell=4$ may not appear to be a large enough value to justify the eikonal approximation, even the $\ell=4$ case provides very accurate predictions, as has already been pointed out by the existing literature [19, 15].
In particular for this case, $\rho+P_{\theta}=-\frac{9\,k\,w_{q}(1+w_{q})}{r^{3(1+w_{q})}}$, then the energy conditions reduce to
NEC: $k\,w_{q}(1+w_{q})\leq 0\;\;\overset{\circ}{\Longrightarrow}\;\;\begin{cases}k>0:&w_{q}\in\left[-\tfrac{1}{3},\,0\right],\\[2.0pt] k<0:&w_{q}\notin\left(-\tfrac{1}{3},\,0\right),\end{cases}$ (3.25)
[…]
SEC: $k\,w_{q}(1+w_{q})\leq 0\;\;\land\;\;k\,w_{q}(1+3w_{q})\leq 0\;\;\overset{\circ}{\Longrightarrow}\;\;\begin{cases}k>0:&w_{q}\in\left[-\tfrac{1}{3},\,0\right],\\[2.0pt] k<0:&w_{q}\notin\left(-1,\,0\right),\end{cases}$ (3.27)
DEC: $\bigl|k\,w_{q}(1+3w_{q})\bigr|\leq-2k\,w_{q}\;\;\overset{\circ}{\Longrightarrow}\;\;\begin{cases}k>0:&w_{q}\in\left[-1,\,0\right],\\[2.0pt] k<0:&w_{q}\in\left(0,\,\tfrac{1}{3}\right),\end{cases}$ (3.28)
Unlike the previous cases, the energy conditions are not dependent on the position of the UCOP radii. All conditions are compatible with a small $|k|$.
In figures figure 5(a) and 5(b), we demonstrate the damping oscillation waveforms $\log{(|\Psi|)}$ for Schwarzschild, Bardeen, Hayward, and Kiselev models. One can clearly see the decay of the amplitude in time, exhibiting the damped oscillation. The slope connecting the maximal points in the waveforms describes the damping rate, while the spacing between the cusps represents the oscillation frequency. Across all models, the QNM amplitude exhibits the expected damped ringdown, but the details of the decay differ noticeably between cases. In figure 5(a), we can observe how, in comparison to the vacuum scenario, signals for Bardeen and Hayward travel with a similar decaying rate, but with a clearly different frequency. In figure 5(b), we can observe different signals for the Kiselev model according to the $w_{q}$ value. There is a clear difference in the slope connecting the maximal points. In particular, the case $w_{q}=-1$ has a gentle downward slope, therefore a smaller decay rate. On the other hand the cusps are more distanced from each other in comparison with the other cases. It represents a slower oscillation with a slower decay.
4 Stationary rotating hairy black hole and QNMs
In this section, we consider stationary rotating hairy black holes. For rotating black holes, the correspondence between the UCOP and QNMs does not necessarily hold in equal footing as for the static case. Rotation breaks the spherical symmetry, therefore only the equatorial photon orbits control the eikonal QNMs. Consequently, the correspondence applies specifically for the modes with $\ell=|\mathbf{m}|$, which corresponds to UCOP on the equatorial plane [14, 15, 16]. The co-rotating and counter-rotating rays correspond to $\mathbf{m}=\mp\ell$. In this limit, the QNM frequencies take the form
$\omega_{QNM}=\mathbf{m}\,\Omega_{\varepsilon}-i\,(n+1/2)\,\lambda_{\varepsilon}\,$, (4.1)
where we introduce the label $\varepsilon=+1$ for the counter-rotating and $\varepsilon=-1$ for co-rotating equatorial photon orbits. Our sign conventions are summarized in Table1.
Table 1: Sign and range conventions for the rotating-case analysis.
$\varepsilon$ $a$ $\Omega_{\varepsilon}$ $\lambda_{\varepsilon}$ $\delta{\Omega_{\varepsilon}}$ $\delta{\hat{\Omega}_{\varepsilon}}$ $\delta{\lambda_{\varepsilon}}$
Co-rotating - ++ ++ ++ ++ ++ $\pm$
Counter-rotating ++ ++ - ++ - ++ $\pm$
4.1 Rotating Hairy Black Hole and null geodesics
We focus on the following form of the metric
$ds^{2}$ $=$ $-fdt^{2}-2a\sin^{2}\theta(1-f)dtd\varphi+\left\{\Sigma+(2-f)a^{2}\sin^{2}\theta\right\}\sin^{2}\theta d\varphi^{2}$ (4.2)
$\qquad+\dfrac{\Sigma}{\Sigma f+a^{2}\sin^{2}\theta}dr^{2}+\Sigma d\theta^{2}\,$,
$f:=1-\dfrac{2m(r)r}{\Sigma}\,,\quad\Sigma:=r^{2}+a^{2}\cos^{2}\theta\,$,
where $m(r)$ is an arbitrary differentiable function of $r$, and $a$ denotes the angular momentum parameter. For $m(r)=M$, this metric reduces to the Kerr solution. The above metric can be obtained from the static metric (A.1) by setting $f(r)=1/h(r)$ and performing Janis-Newman transformation [17, 18]. This class of metric encompasses a fairly large class of stationary rotating hairy black holes, including the rotating Bardeen, the rotating Hayward, the rotating Kiselev, the quantum improved Kerr [22], among others. This metric can be defined by replacing the mass function $m(r)$ by its corresponding value in the static case, defined as in expression (2.29).
Let us consider circular geodesic orbits. We restrict our attention to the equatorial plane $\theta=\pi/2$. The Lagrangian is given as
${\cal L}=\dfrac{1}{2}\left(-f\dot{t}^{2}+\dfrac{r^{2}\dot{r}^{2}}{a^{2}+r^{2}f}-2a(1-f)\dot{t}\dot{\varphi}+[r^{2}+(2-f)a^{2}]\dot{\varphi}^{2}\right)=\dfrac{\epsilon}{2}\,$, (4.3)
where $\epsilon=1$ for the timelike case or $\epsilon=0$ for the null case. As for the static case, we define the two conserved quantities $E$ and $L$ as
$E:=-\dfrac{\partial{\cal L}}{\partial\dot{t}}=f\dot{t}+a(1-f)\dot{\varphi}\,,\quad L:=\dfrac{\partial{\cal L}}{\partial\dot{\varphi}}=-a(1-f)\dot{t}+[r^{2}+(2-f)a^{2}]\dot{\varphi}\,.$ (4.4)
Then, by solving for $\dot{t}$ and $\dot{\varphi}$, we obtain
$\dot{t}=\dfrac{1}{\triangle}\left\{\left(r^{2}+a^{2}+\dfrac{2m(r)a^{2}}{r}\right)E-\dfrac{2m(r)a}{r}L\right\}\,,\quad\dot{\varphi}=\dfrac{1}{\triangle}\left\{\dfrac{2m(r)a}{r}E+\left(1-\dfrac{2m(r)}{r}\right)L\right\}\,$, (4.5)
where $\triangle:=a^{2}+r^{2}f=r^{2}+a^{2}-2m(r)r$. Then, from the expression of the Lagrangian, we find
$\dfrac{1}{2}\dot{r}^{2}+V(r)=0\,,\quad V(r):=\dfrac{1}{2r^{2}}\left[L^{2}-(r^{2}+a^{2})E^{2}-\dfrac{2m(r)}{r}(aE-L)^{2}-\triangle\epsilon\right]\,.$ (4.6)
The conditions for circular orbits with radius $r_{\star}$ are $V(r_{\star})=0$ and $V^{\prime}(r_{\star})=0$. From these conditions, we have
$(a^{2}+r^{2})E^{2}-L^{2}+\dfrac{2m}{r}(aE-L)^{2}+\triangle\epsilon=0\,,\quad 2E^{2}r+\left(\dfrac{2m^{\prime}}{r}-\dfrac{2m}{r^{2}}\right)(aE-L)^{2}+\triangle^{\prime}\epsilon=0\,.$ (4.7)
For the null geodesic case ($\epsilon=0$), it is convenient to introduce the impact parameter $D:=L/E$. Then, the two circular-orbit conditions become
$a^{2}-D^{2}+r^{2}+\dfrac{2m}{r}(a-D)^{2}\circeq 0\,,\quad 2r^{2}+\left(2m^{\prime}-\dfrac{2m}{r}\right)(a-D)^{2}\circeq 0\,$, (4.8)
and we find that the impact parameter and the UCOP radii should satisfy
$D_{\varepsilon}\circeq a-\varepsilon\sqrt{\dfrac{r^{3}}{m-rm^{\prime}}}\,$, (4.9)
$r\circeq(3m(r)-rm^{\prime}(r))\dfrac{D_{\varepsilon}-a}{D_{\varepsilon}+a}.$ (4.10)
The second derivative of the potential is given by
$V^{\prime\prime}\circeq-\dfrac{E^{2}(D_{\varepsilon}-a)^{2}}{r^{5}}\left[3(m-rm^{\prime})+r^{2}m^{\prime\prime}\right]\,.$ (4.11)
From the standard definitions of the orbital frequency and the Lyapunov exponent, we derive
$\Omega_{\varepsilon}\circeq\dfrac{1}{D_{\varepsilon}}\,,\quad\lambda_{\varepsilon}^{2}\circeq\left(\dfrac{D_{\varepsilon}+a}{D_{\varepsilon}}\right)^{2}\dfrac{\triangle^{2}}{r^{3}}\dfrac{3(m-rm^{\prime})+r^{2}m^{\prime\prime}}{[(D_{\varepsilon}-2a)m-D_{\varepsilon}\\ \ r\\ \ m^{\prime}]^{2}}.$ (4.12)
Note that for the non-rotating vacuum case, i.e., $a=0$, $m=M$, the above formula reduces to the Schwarzschild result $r_{\star,\varepsilon}=r_{0}=3M$.
Now by using the Einstein equations, we find (in Appendix B) that
$m^{\prime}=-\dfrac{r^{2}}{2}G^{r}{}_{r}=-4\pi r^{2}P_{r}\,,\quad m^{\prime\prime}=-rG^{\theta}{}_{\theta}=-8\pi rP_{\theta}\,.$ (4.13)
Therefore, at the UCOP radius position $r_{\star,\varepsilon}$, we find
$\Omega_{\varepsilon}$ $=D_{\varepsilon}^{-1}=\left(a-\varepsilon\sqrt{\dfrac{{r_{\star,\varepsilon}}^{3}}{m+4\pi{r_{\star,\varepsilon}}^{3}P_{r}}}\right)^{-1}\,$$\quad$ (4.14)
$\lambda_{\varepsilon}$ $=\bigg|\dfrac{D_{\varepsilon}+a}{D_{\varepsilon}}\dfrac{\triangle}{{r_{\star,\varepsilon}}^{3/2}}\dfrac{(3m-4\pi{r_{\star,\varepsilon}}^{3}(2P_{\theta}-3P_{r}))^{1/2}}{(D_{\varepsilon}-2a)m+4\pi D_{\varepsilon}\,{r_{\star,\varepsilon}}^{3}P_{r}}\bigg|\,$, (4.15)
where all quantities are evaluated at $r=r_{\star,\varepsilon}$.
In order to model these general expressions as a small hair perturbation introduced to a vacuum Kerr black hole, we expand the expressions up to the first order
$r_{\star,\varepsilon}\simeq r_{K,\varepsilon}+\delta r_{\varepsilon}\,$$\quad$ (4.16)
$m(r)\simeq M+\delta m(r)\,$$\quad$ (4.17)
$D_{\varepsilon}\simeq D_{K,\varepsilon}+\delta D_{\varepsilon}\,$$\quad$ (4.18)
$\Omega_{\varepsilon}\simeq\Omega_{K,\varepsilon}+\delta\Omega_{\varepsilon}\,$$\quad$ (4.19)
$\lambda_{\varepsilon}\simeq\lambda_{K,\varepsilon}+\delta\lambda_{\varepsilon}\,$$\quad$ (4.20)
where $r_{K,\varepsilon}$, $D_{K,\varepsilon}$, $\Omega_{K,\varepsilon}$, $\lambda_{K,\varepsilon}$ are the values of the parameters in the Kerr case for the branch $\varepsilon$. For the Kerr black hole, the corresponding quantities are given by
$r_{K,\varepsilon}=2M\left(1+\cos{\left(\frac{2}{3}\arccos{\left(-\frac{\varepsilon\,a}{M}\right)}\right)}\right)\,$$\quad$ (4.21)
$D_{K,\varepsilon}=a-\varepsilon\sqrt{\frac{r_{K,\varepsilon}^{3}}{M}}\,$$\quad$ (4.22)
$\Omega_{K,\varepsilon}=\frac{1}{D_{K,\varepsilon}}\,$$\quad$ (4.23)
$\lambda_{K,\varepsilon}=\bigg|\frac{\sqrt{3}\,\triangle(D_{K,\varepsilon}+a)}{\sqrt{M}r_{K,\varepsilon}^{3/2}\,D_{K,\varepsilon}\,(D_{K,\varepsilon}-2a)}\bigg|=\frac{\sqrt{3}(r_{K,\varepsilon}-M)}{r_{K,\varepsilon}(r_{K,\varepsilon}+3M)},$ (4.24)
where the allowed radii are constrained by the extremal Kerr limit $|a|=M$. The co-rotating branch ($\varepsilon=-1$) can take values in the interval $[M,3M]$, while the counter-rotating branch lies in the interval $[3M,4M]$.
For the orbital frequency (4.14), we obtain
$\delta D_{\varepsilon}$ $=\varepsilon\frac{{r_{K,\varepsilon}}^{3/2}}{2\,M^{1/2}}\Big(\frac{\delta m-{r_{K,\varepsilon}}\delta m^{\prime}}{M}-\frac{3\delta r_{\varepsilon}}{r_{K,\varepsilon}}\Big),$ (4.25)
$\delta\Omega_{\varepsilon}$ $=-\,\Omega_{K,\varepsilon}^{2}\,\delta D_{\varepsilon}.$ (4.26)
To calculate the shift of the position of the photon orbits radii we proceed to eliminate $D$ from the null circular-orbit conditions (4.9) and (4.10). Then we obtain the following implicit UCOP radius equation
$F_{\varepsilon}(r;m,m^{\prime},a):=r^{2}\big(3m-r-rm^{\prime}\big)+2\,a\,\varepsilon\,r^{3/2}\sqrt{m-rm^{\prime}}\circeq 0,$ (4.27)
which reduces to $r(3M-r)+2a\,\varepsilon\sqrt{Mr}\circeq 0$ for the Kerr case. Linearizing $F_{\varepsilon}(r_{K,\varepsilon}+\delta r_{\varepsilon},M+\delta m,\delta m^{\prime};a)=0$ gives
$\delta r_{\varepsilon}\circeq-\frac{(\partial F_{\varepsilon}/\partial m)_{K}\,\delta m+(\partial F_{\varepsilon}/\partial m^{\prime})_{K}\,\delta m^{\prime}}{(\partial F_{\varepsilon}/\partial r)_{K}}\Bigg|_{r=r_{\mathrm{K,\varepsilon}}},$ (4.28)
where the subscript $K$ indicates that the corresponding quantities are evaluated in the Kerr spacetime, i.e., at $r=r_{K,\varepsilon}$ with $m(r)=M$,
$\left(\frac{\partial F_{\varepsilon}}{\partial r}\right)_{K}=\frac{3}{2}\,r_{K,\varepsilon}(M-r_{K,\varepsilon}),\quad\left(\frac{\partial F_{\varepsilon}}{\partial m}\right)_{K}=\frac{{r_{K,\varepsilon}}^{2}(r_{K,\varepsilon}+3\,M)}{2M},\quad\left(\frac{\partial F_{\varepsilon}}{\partial m^{\prime}}\right)_{K}=-\,\frac{{r_{K,\varepsilon}}^{3}(r_{K,\varepsilon}-M)}{2M}.$ (4.29)
Then the perturbation on the photon orbit positions is given by
$\delta r_{\varepsilon}\simeq-\frac{1}{3M}\left({r_{K,\varepsilon}}^{2}\,\delta m^{\prime}-\frac{{r_{K,\varepsilon}}({r_{K,\varepsilon}}+3\,M)}{{r_{K,\varepsilon}}-M}\,\delta m\right)\Bigg|_{r=r_{K,\varepsilon}}.$ (4.30)
Substituting the expression for $\delta r_{\varepsilon}$ (4.30) into the expression for $\delta D_{\varepsilon}$ (4.26), we obtain
$\delta D_{\varepsilon}\simeq-\frac{2\,\varepsilon}{r_{K,\varepsilon}-M}\sqrt{\frac{{r_{K,\varepsilon}}^{3}}{M}}\,\delta m\Bigg|_{r=r_{K,\varepsilon}},$ (4.31)
which gives the perturbation of the impact parameter for the UCOPs. The corresponding orbital frequency is then expressed as
$\delta\Omega_{\varepsilon}\simeq\varepsilon\,D_{K,\varepsilon}^{-2}\left(\frac{2}{r_{K,\varepsilon}-M}\sqrt{\frac{{r_{K,\varepsilon}}^{3}}{M}}\right)\delta m\Bigg|_{r=r_{K,\varepsilon}}\quad.$ (4.32)
By expanding up to the first order, we obtain the Lyapunov exponent as follows
$\delta(\lambda_{\varepsilon}^{2})\simeq\frac{1}{M\,r_{K,\varepsilon}^{2}(3M+r_{K,\varepsilon})^{3}}\left(-2(r_{K,\varepsilon}^{3}+13M\,r_{K,\varepsilon}^{2}-21M^{2}\,r_{K,\varepsilon}-9M^{3})\delta m\right.\\ \left.+2\,r_{K,\varepsilon}(r_{K,\varepsilon}-M)(r_{K,\varepsilon}+3M)(r_{K,\varepsilon}-5M)\delta m^{\prime}+r_{K,\varepsilon}^{2}(r_{K,\varepsilon}-M)^{2}(r_{K,\varepsilon}+3M)\delta m^{\prime\prime}\right)\Bigg|_{r=r_{K,\varepsilon}}.$ (4.33)
Finally we obtain the perturbation by using $\delta\lambda=(2\lambda_{K,\varepsilon})^{-1}\delta(\lambda^{2})$ as
$\delta\lambda_{\varepsilon}\simeq\frac{1}{2\sqrt{3}M(3M+r_{K,\varepsilon})}\Bigg(\frac{2(9M^{3}+21M^{2}\,r_{K,\varepsilon}-13M\,r_{K,\varepsilon}^{2}-r_{K,\varepsilon}^{3})}{r_{K,\varepsilon}(r_{K,\varepsilon}-M)(3M+r_{K,\varepsilon})}\delta m\\ +2(r_{K,\varepsilon}-5M)\delta m^{\prime}+(r_{K,\varepsilon}-M)r_{K,\varepsilon}\,\delta m^{\prime\prime}\Bigg)\Bigg|_{r=r_{K,\varepsilon}}.$ (4.34)
For $a=0$, from the expression (4.32), we obtain
$\delta\Omega_{\varepsilon}\simeq\varepsilon\left(\frac{M}{r_{K,\varepsilon}^{3}}\right)\left(\frac{2}{r_{K,\varepsilon}-M}\sqrt{\frac{r_{K,\varepsilon}^{3}}{M}}\right)\delta m\\ =\frac{2\,\varepsilon\sqrt{M}}{r_{K,\varepsilon}^{3/2}(r_{K,\varepsilon}-M)}\,\delta m=\frac{\varepsilon}{3^{3/2}M^{2}}\,\delta m,$ (4.35)
where we have used $r_{K,\varepsilon}=3M$ for $a=0$. By substituing $\delta m=-\frac{r}{2}\,\delta f$ into the previous result,
$\delta\Omega_{\varepsilon}\simeq\varepsilon\,\frac{1}{3^{3/2}M^{2}}\left(-\frac{3M}{2}\,\delta f\right)=-\varepsilon\,\frac{1}{2\sqrt{3}\,M}\,\delta f.$ (4.36)
By picking the branch $\varepsilon=-1$, or computing $\delta\hat{\Omega}_{\varepsilon}$, we recover the static result.
In a similar way, the shift of the Lyapunov exponent for $a=0$ can be evaluated as
$\delta\lambda_{\varepsilon}\simeq\frac{1}{12\sqrt{3}\,M^{2}}\left(6M^{2}\,\delta m^{\prime\prime}-4\,\delta m-4M\,\delta m^{\prime}\right)\Bigg|_{r=3M}.$ (4.37)
Rewriting the mass perturbations using $\delta m=-\tfrac{r}{2}\,\delta f$ and substituting the Einstein equations (B.10), we obtain
$\delta\lambda_{\varepsilon}\simeq\left[\frac{1}{2\sqrt{3}\,M}\,\delta f(3M)-4\sqrt{3}\,\pi M\bigl(P_{\theta}(3M)-P_{r}(3M)\bigr)\right].$ (4.38)
It reproduces the static Lyapunov exponent shift.
4.2 Modifications of QNMs and Energy Conditions
We continue the analysis from the static case. Although an analytically simple expression for $\lambda_{\varepsilon}$ is not available, as for the static case, we can quantify the expression in terms of the hair characteristics.
By applying the Einstein equations in the rotating metric form (4.2), as explained in the appendix B, in particular for the equatorial plane (B.10), we can express the derivatives of $m(r)$ as
$m^{\prime}(r)=\delta m^{\prime}(r)=-4\pi r^{2}P_{r}(r),$ (4.39)
$m^{\prime\prime}(r)=\delta m^{\prime\prime}(r)=-8\pi rP_{\theta}(r).$ (4.40)
We start from the Einstein equation (4.39) for the deviation of the mass function. We impose the boundary condition that the deviation vanishes at infinity $\delta m(\infty)=0$. Integrating $\delta m^{\prime}(r)$ from $r$ to $\infty$ gives
$\delta m(r)=-\int_{r}^{\infty}\delta m^{\prime}(s)\,ds\,=4\pi\int_{r}^{\infty}s^{2}P_{r}(s)\,ds.$ (4.41)
For simplicity, we use (4.41) to express the components of the QNM as
$\delta\hat{\Omega}_{\varepsilon}$ $=$ $-\varepsilon\,\delta\Omega_{\varepsilon}\simeq-D_{K,\varepsilon}^{-2}\left(\frac{8\pi}{r_{K,\varepsilon}-M}\sqrt{\frac{r_{K,\varepsilon}^{3}}{M}}\right)\int_{r_{K,\varepsilon}}^{\infty}s^{2}P_{r}(s)\,ds\,$, (4.42)
$\delta\lambda_{\varepsilon}$ $\simeq$ $\,\frac{2\pi}{\sqrt{3}M\,\bigl(3M+r_{K,\varepsilon}\bigr)}\Bigg[A(r_{K,\varepsilon})\int_{r_{K,\varepsilon}}^{\infty}s^{2}P_{r}(s)\,ds$ (4.44)
$\hskip 56.9055pt+B(r_{K,\varepsilon})\,r_{K,\varepsilon}^{2}P_{r}+2\,C(r_{K,\varepsilon})\,r_{K,\varepsilon}P_{\theta}\Bigg],$
where
$A(r)\equiv\frac{2\left(r^{3}+13Mr^{2}-21M^{2}r-9M^{3}\right)}{r\,(r-M)\,(3M+r)}\,,\qquad B(r)\equiv 2(r-5M)\,,\qquad C(r)\equiv(r-M)\,r\,.$ (4.45)
The sign of the shift of the UCOP radii $\delta r_{\varepsilon}$ (4.30) is not so clear as in the static case, while the term that contains $\delta m$ is a positive contribution, the other term is proportional to $P_{r}$, and there is no restriction on the sign of this term. The shift of the UCOP radii $\delta r_{\varepsilon}$ can be expressed as
$\delta r_{\varepsilon}\simeq\frac{4\pi}{3M}\,r_{K,\varepsilon}^{4}\,P_{r}(r_{K,\varepsilon})+\frac{4\pi}{3M}\frac{r_{K,\varepsilon}(r_{K,\varepsilon}+3M)}{r_{K,\varepsilon}-M}\,\int_{r_{K,\varepsilon}}^{\infty}s^{2}P_{r}(s)\,ds.$ (4.46)
In Appendix B, we derived the values for the energy-momentum tensor components at the co-rotating frame in our settings. The values for the energy density and the pressures at the co-rotating frame are the same values as in the static case. Therefore, the analysis of the energy conditions (2.35), (2.36), (2.37), and (2.38) remains valid. Energy conditions are frame-independent, therefore the interpretation in the co-rotating frame should hold true for any frame. Furhermore, by following the definitions (2.15), from expressions listed in Appendix B, we can indentify $w_{r}=-1$, and
$w_{\theta}=-\frac{r}{2}\frac{m^{\prime\prime}}{m^{\prime}}\,.$ (4.47)
Then, according to (4.42) and  (4.46), $\delta\hat{\Omega}_{\varepsilon}$ must be positive if we consider a positive energy density $\rho$, as seen in the appendix B, and $\delta r_{\varepsilon}$, negative, as in the static case.
5 Examples of Stationary Rotating Hairly Black Holes
5.1 Rotating Bardeen Black Hole
The rotating Bardeen geometry generated via the Newman-Janis prescription can be described by taking the same mass function $m(r)=M\cdot[r^{2}/(r^{2}+q^{2})]^{3/2}$ as in the static solution. In the small-$q$ regime, one finds the following expansion:
$m(r)\;=\;M\Big(1-\frac{3\,q^{2}}{2r^{2}}\Big)\;+\;\mathcal{O}(q^{4}),\qquad m^{\prime}(r)\;\simeq\;\,\frac{3Mq^{2}}{r^{3}},\qquad m^{\prime\prime}(r)\;\simeq\;-\,\frac{9Mq^{2}}{r^{4}}.$ (5.1)
On the equatorial plane ($\theta=\pi/2$), the Einstein equations for the stationary axisymmetric ansatz (B.5), (B.6), (B.11) and (B.12) relate the derivatives of $m(r)$ to the effective anisotropic pressures. Keeping only the leading contribution in $q$, we obtain
$P_{r}(r)\simeq-\,\frac{3Mq^{2}}{4\pi r^{5}},\qquad P_{\theta}(r)\simeq\,\frac{9Mq^{2}}{8\pi r^{5}}.$ (5.2)
We now insert the perturbative expansion of $m(r)$ into the impact-parameter relation (4.9) and expand consistently to first order in the deformation parameter (i.e. $\mathcal{O}(q^{2})$). This yields the shifts in the Kerr photon-sphere radius and in the impact parameter
$\delta r_{\varepsilon}$ $\simeq-\,\frac{3r_{K,\varepsilon}+M}{2r_{K,\varepsilon}(r_{K,\varepsilon}-M)}\,q^{2}\,$, (5.3)
$\delta D_{\varepsilon}$ $\simeq\frac{3\,\varepsilon\,q^{2}}{r_{K,\varepsilon}-M}\sqrt{\frac{M}{r_{K,\varepsilon}}}\,.$ (5.4)
We can observe in Figure 6(a) how the rotation splits the UCOP in two, and that as the rotation frequency $a$ increases, the separation between both radii increases. Given the presence of $q^{2}$ in (5.3), we can observe that for any value of $q$, the radii of the UCOP reduce.
(a) UCOP radius change given different $q$ parameters for different angular momentum values. The dashed lines show the cases of the Kerr solutions, and the solid lines represent the cases of the rotating Bardeen Black Hole.
(b) Deviation of the angular frequency $\delta\hat{\Omega}_{\varepsilon}$ as a function of $q/M$ for several values of the spin parameter $a$. Solid lines and dashed lines represent the counter-rotating and co-rotating solution branches, respectively.
(c) Deviation of the Lyapnov exponent $\delta\lambda_{\varepsilon}$ as a function of $q/M$ for several values of the spin parameter $a$. Solid lines and dashed lines represent the counter-rotating and co-rotating solution branches, respectively.
Figure 6: Plots for rotating Bardeen black holes
The corresponding corrections to the angular frequency and Lyapunov exponent take the forms
$\delta\hat{\Omega}_{\varepsilon}=-\varepsilon\delta\Omega_{\varepsilon}\simeq\left(a-\varepsilon\sqrt{\frac{{r_{K,\varepsilon}}^{3}}{M}}\right)^{-2}\left(\frac{3\,q^{2}}{r_{K,\varepsilon}-M}\sqrt{\frac{M}{r_{K,\varepsilon}}}\right),$ (5.5)
$\delta\lambda_{\varepsilon}\simeq\frac{2\,\sqrt{3}M\,q^{2}}{(r_{K,\varepsilon}-M)\,r_{K,\varepsilon}^{3}\,(r_{K,\varepsilon}+3M)^{2}}\left(3M^{2}-8Mr_{K,\varepsilon}+r_{K,\varepsilon}^{2}\right).$ (5.6)
Figure 6(b) displays the deviation in the angular frequency $\delta\hat{\Omega}_{\varepsilon}$ of the unstable circular photon orbit as a function of the parameter $q/M$. The deviation is positive and grows monotonically with $q/M$. The yellow line in figure 6(b) represents the case when the rotating cases reduce to the static case by setting the spin $a$ to 0. Both counter-rotating and co-rotating cases converge to a single curve, as expected. We can see that for the counter-rotating case, the magnitude of the shift produced by the hair is larger. The opposite is true for the co-rotating case. As expected, all curves converge to $\delta\hat{\Omega}_{\varepsilon}=0$ in the Kerr limit $q/M\to 0$.
Figure 6(c) shows the behaviour of the deviation parameter $\delta\lambda_{\varepsilon}$ as a function of $q/M$ for several values of the spin parameter $a$. As in figure 6(c), the yellow line represents the static case. The magnitude of the shift for the counter-rotating case is larger in magnitude, and that of the co-rotating case is smaller. All curves smoothly approach $\delta\lambda_{\varepsilon}=0$ in the limit $q/M\to 0$, confirming that the Kerr solution is recovered continuously.
5.2 Rotating Hayward Black Hole
In the rotating Hayward black hole case, the mass function is defined as $m(r)=Mr^{3}/(r^{3}+q^{3})$, and for small $q$ it admits the expansion
$m(r)\;=\;M\Big(1-\frac{\,q^{3}}{r^{3}}\Big)\;+\;\mathcal{O}(q^{4}),\qquad m^{\prime}(r)\;\simeq\;\,\frac{3Mq^{3}}{r^{4}},\qquad m^{\prime\prime}(r)\;\simeq\;-\,\frac{12Mq^{3}}{r^{5}}.$ (5.7)
Restricting to the equatorial plane and using the same set of field equations, we obtain, at leading order in $q$,
$P_{r}(r)\simeq-\,\frac{3Mq^{3}}{4\pi r^{6}},\qquad P_{\theta}(r)\simeq\,\frac{3Mq^{3}}{2\pi r^{6}}.$ (5.8)
Proceeding as in the Bardeen case, we expand (4.9) consistently to the first non-vanishing order. This gives
$\delta r_{\varepsilon}$ $\simeq-\,\frac{4\,q^{3}}{3r_{K,\varepsilon}(r_{K,\varepsilon}-M)}\,$, (5.9)
$\delta D_{\varepsilon}$ $\simeq\frac{2\,\varepsilon\,q^{3}}{r_{K,\varepsilon}-M}\sqrt{\frac{M}{r_{K,\varepsilon}^{3}}}\,.$ (5.10)
The induced corrections to the angular frequency and the Lyapunov exponent are therefore
$\delta\hat{\Omega}_{\varepsilon}\simeq\left(a-\varepsilon\sqrt{\frac{{r_{K,\varepsilon}}^{3}}{M}}\right)^{-2}\left(\frac{2\,q^{3}}{r_{K,\varepsilon}-M}\sqrt{\frac{M}{{r_{K,\varepsilon}}^{3}}}\right),$ (5.11)
$\delta\lambda_{\varepsilon}\simeq\frac{2\,q^{3}}{\sqrt{3}\,(r_{K,\varepsilon}-M)\,r_{K,\varepsilon}^{4}\,(r_{K,\varepsilon}+3M)^{2}}\left(9M^{3}-15M^{2}r_{K,\varepsilon}-Mr_{K,\varepsilon}^{2}-r_{K,\varepsilon}^{3}\right).$ (5.12)
The set of plots in figure 7 exhibits the same qualitative behaviour as the Bardeen case. The deviations remain smooth, monotonic as functions of in the parameter $q/M$, and approach the Kerr limit continuously as $q/M\to 0$, with only a mild dependence on the spin parameter $a$.
(a) UCOP radius change given different $q$ parameters for different angular momentum values. The dashed lines show the cases of the Kerr solutions, and the solid lines represent the cases of the rotating Hayward Black Hole.
(b) Deviation of the angular frequency $\delta\hat{\Omega}_{\varepsilon}$ as a function of $q/M$ for several values of the spin parameter $a$.
(c) Deviation of the Lyapnov exponent $\delta\lambda_{\varepsilon}$ as a function of $q/M$ for several values of the spin parameter $a$.
Figure 7: Plots for rotating Hayward black holes
5.3 Rotating Kiselev Black Hole
For the rotating Kiselev solution the mass function is
$m(r)=M+\frac{k}{2r^{3\,w_{q}}}\;,\quad m^{\prime}(r)=-\,\frac{3k\,w_{q}}{2r^{3w_{q}+1}},\quad m^{\prime\prime}(r)=\frac{3k\,w_{q}(1+3w_{q})}{2r^{3w_{q}+2}}.$ (5.13)
On the equatorial plane, the Einstein equations yield the effective pressures
$P_{r}(r)=\frac{3k\,w_{q}}{8\pi r^{3(1+w_{q})}},\qquad P_{\theta}(r)=-\frac{3k\,w_{q}\,(1+3w_{q})}{16\pi\,r^{3(1+w_{q})}}.$ (5.14)
Expanding the impact parameter (4.9) to first order in the deformation (i.e. linear in $k$) we obtain
$\delta r_{\varepsilon}$ $\simeq\frac{k\,\left(r_{K,\varepsilon}\right)^{1-3w_{q}}\,\Bigl(r_{K,\varepsilon}-3M(-1+w_{q})+3r_{K,\varepsilon}w_{q}\Bigr)}{6M\,(r_{K,\varepsilon}-M)}\,$, (5.15)
$\delta D_{\varepsilon}$ $\simeq\frac{\varepsilon\,k\,\left(r_{K,\varepsilon}\right)^{-3w_{q}}\,\sqrt{(r_{K,\varepsilon})^{3}/M}}{r_{K,\varepsilon}-M}\,.$ (5.16)
Accordingly, the first-order corrections to the angular frequency and the Lyapunov exponent read
$\delta\hat{\Omega}_{\varepsilon}\simeq\left(a-\varepsilon\sqrt{\frac{{r_{K,\varepsilon}}^{3}}{M}}\right)^{-2}\left(\frac{-\,k\,(r_{K,\varepsilon})^{-3\,w_{q}}\,\sqrt{(r_{K,\varepsilon})^{3}/M}}{r_{K,\varepsilon}-M}\right),$ (5.17)
$\delta\lambda_{\varepsilon}\simeq\frac{-\,k\,r_{K,\varepsilon}^{-1-3w_{q}}}{4\sqrt{3}\,M\,(M-r_{K,\varepsilon})\,(3M+r_{K,\varepsilon})^{2}}\left(9M^{3}\!\left(2+3(-3+w_{q})w_{q}\right)\right.\\ \left.+3M^{2}\!\left(14+3(7-5w_{q})w_{q}\right)r_{K,\varepsilon}+M\!\left(-26+3w_{q}(7+3w_{q})\right)r_{K,\varepsilon}^{2}\right.\\ \left.+(-2+3w_{q})(1+3w_{q})\,r_{K,\varepsilon}^{3}\right).$ (5.18)
(a) UCOP radius change given different $w_{q}$ parameters. Blue represents the counter-rotating branch solution and orange the co-rotating branch solution. The dotted lines show the cases of the Kerr solutions, and the dashed lines represent the cases of the rotating Kiselev black hole model.
(b) Deviation of the angular frequency $\delta\hat{\Omega}_{\varepsilon}$ as a function of $w_{q}$ for several values of the spin parameter $a$ with k=0.04k=-0.04.
(c) Deviation of the Lyapnov exponent $\delta\lambda_{\varepsilon}$ as a function of $w_{q}$ for several values of the spin parameter $a$ with k=0.04k=-0.04.
Figure 8: Plots for rotating Kiselev black holes
The behaviour of the QNM coefficients for the Kiselev black hole, shown in Figs. 8, can be interpreted directly from the effective matter content of the model. The parameters $k$ and $w_{q}$ govern the effective energy-momentum distribution, which in turn determines the sign and radial profile of the deviation $\delta m(r)$ and the anisotropic pressures $(P_{r},P_{\theta})$.
When $w_{q}$ is in the range of quintessence, the shift for the Lyapunov exponent is positive. Unlike the Bardeen and Hayward cases, where the matter distribution tended to reduce the instability, here the effective matter contribution enhances the divergence rate of nearby null geodesics. Outside of this range, we can find values of $w_{q}$ that produce the same sign values of $\delta\Omega$ and $\delta\lambda$ as in the previous cases. This is clearly visible in the plots 8(b) and 8(c).
6 Summary and discussions
In this paper, we derived the expressions of the QNM frequencies for stationary hairy black holes by using the correspondence between the circular null geodesics and QNMs. We modeled the deviation from the corresponding vacuum solution for the Einstein equations as perturbations in the form of an anisotropic fluid. Then, we utilized the relation between the eikonal limit of QNMs and null circular orbits. We explicitly obtained QNM formulas expressed in terms of the state parameters of the hair field. Our formulas provide a systematic approach to the analysis of QNMs for hairy black holes without assuming any specific underlying theory or model. We also examined the relationship between the state parameters and the QNM frequencies, including possible violations of the energy conditions. For the first two static models discussed in this work, the small-deviation approximation from the vacuum solution is incompatible with the dominant energy condition (DEC). This implies that, if the corresponding deviation is confirmed through the QNM observations, we need to consider a matter field violating the DEC or any other modification in the gravitational wave emission mechanisms. Since the violation of the DEC implies a superluminal flux, it might be healthier to consider a modified gravity theory than to rely on an exotic matter field. WEC and SEC are satisfied for all our examples, therefore solutions still maintain a positive energy density, and gravity is still locally attractive in both timelike and null geodesics.
In our model, we considered the rotating Bardeen, Hayward, and Kiselev metric constructed by applying Janis-Newman method to their static counterpart. Then we restricted our attention to the UCOP on the equatorial plane and obtain angular frequency and Lyapunov exponent to get the QNM information. The correspondence between the QNM and the UCOP for the rotating case is not yet fully established, but for the UCOP on the equatorial plane its considered to be related to the $l=|m|$ mode of QNM [2].
In the construction of our hairy black hole models, we considered the effects of matter fields as perturbative deviations from the Schwarzschild and the Kerr solutions. In particular, we assumed the simplest form of the equations of state (2.15) and focused only on the first order perturbative effects of the matter field. It should be noted that our formulas given in Sec.  2.1 (also the formula (2.14)) for the static case, and those in Sec. 4.1 (see e.g., (4.14), (4.15)) for the rotating case can apply for the general cases. It would therefore be interesting to extend our present analyses to higher order perturbative analyses, just like higher order WKB analyses, and to include more general equation of state, such as the polytrope type. Possible modifications of QNMs induced by matter fields or deviations from general relativity were also discussed by Cardoso et al. [23] and McManus et al. [24], who considered perturbative deformations of the effective potential in the wave equations. It would be interesting to compare their results with those obtained in our analysis.
Acknowledgments
This work was supported in part by JSPS KAKENHI Grant No. JP24K07027 (C.Y.), JP25K07281 (C.Y.), JP25K07306 (A.I.), and also supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A Extreme Universe No. JP21H05182(A.I.) and JP21H05186(A.I.).
Appendix A Einstein tensor for static spherically symmetric metric
Let us consider spherically symmetric, static metric of the form
$ds^{2}=-f(r)dt^{2}+h(r)dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2})\,.$ (A.1)
The non-vanishing components of the connection coeffecients are
${\Gamma}^{r}_{tt}=\dfrac{f^{\prime}}{2h}\,,\quad{\Gamma}^{r}_{rr}=\dfrac{h^{\prime}}{2h}\,,\quad{\Gamma}^{r}_{\theta\theta}=-\dfrac{r}{h}\,,\quad{\Gamma}^{r}_{\varphi\varphi}={\Gamma}^{r}_{\theta\theta}\sin^{2}\theta\,.$ (A.2)
The Einstein tensor is given by the components
$G^{t}{}_{t}$ $=$ $-\dfrac{1}{rh}\left[\dfrac{h^{\prime}}{h}+\dfrac{1}{r}(h-1)\right]\,$, (A.3)
$G^{r}{}_{r}$ $=$ $-\dfrac{1}{rh}\left[-\dfrac{f^{\prime}}{f}+\dfrac{1}{r}(h-1)\right]\,$, (A.4)
$G^{\theta}{}_{\theta}=G^{\varphi}{}_{\varphi}$ $=$ $\dfrac{1}{2\sqrt{fh}}\dfrac{d}{dr}\left(\dfrac{f^{\prime}}{\sqrt{fh}}\right)+\dfrac{1}{2rh}\left(\dfrac{f^{\prime}}{f}-\dfrac{h^{\prime}}{h}\right)$ (A.5)
$=$ $\dfrac{1}{rh}\left[\dfrac{r}{2}\dfrac{f^{\prime\prime}}{f}-\dfrac{r}{4}\left(\dfrac{f^{\prime}}{f}\right)^{2}-\dfrac{4}{r}\dfrac{f^{\prime}}{f}\dfrac{h^{\prime}}{h}+\dfrac{1}{2}\left(\dfrac{f^{\prime}}{f}-\dfrac{h^{\prime}}{h}\right)\right]\,$,
[…]
Appendix B Einstein tensor for a class of stationary axisymmetric metric
Let us consider the following form of stationary axisymmetric metric (4.2)
$ds^{2}$ $=$ $-fdt^{2}-2a\sin^{2}\theta(1-f)dtd\varphi+\left\{\Sigma+(2-f)a^{2}\sin^{2}\theta\right\}\sin^{2}\theta d\varphi^{2}$ (B.1)
$\qquad+\dfrac{\Sigma}{\Sigma f+a^{2}\sin^{2}\theta}dr^{2}+\Sigma d\theta^{2}\,$,
$f:=1-\dfrac{2m(r)r}{\Sigma}\,,\quad\Sigma:=r^{2}+a^{2}\cos^{2}\theta\,$,
with $m(r)$ being an arbitrary (differenciable) function of $r$ and $a$ the spin parameter. The Einstein tensor for the metric (B.1) is given by the components
$G^{t}{}_{t}$ $=$ $\dfrac{1}{\Sigma^{3}}\left\{a^{2}r\sin^{2}\theta\Sigma m^{\prime\prime}+2m^{\prime}[a^{4}\cos^{2}\theta\sin^{2}\theta-a^{2}r^{2}-r^{4}]\right\}\,$, (B.2)
$G^{t}{}_{\varphi}$ $=$ $-\dfrac{1}{\Sigma^{3}}a\sin^{2}\theta\left\{r(a^{2}+r^{2})\Sigma m^{\prime\prime}+2m^{\prime}[a^{4}\cos^{2}\theta-a^{2}r^{2}\sin^{2}\theta-r^{4}]\right\}\,$, (B.3)
$G^{\varphi}{}_{\varphi}$ $=$ $-\dfrac{1}{\Sigma^{3}}\left\{r(a^{2}+r^{2})\Sigma m^{\prime\prime}+2a^{2}m^{\prime}[\Sigma-2r^{2}\sin^{2}\theta]\right\}\,$, (B.4)
$G^{\theta}{}_{\theta}$ $=$ $-\frac{1}{\Sigma^{2}}\left\{r\Sigma m^{\prime\prime}+2a^{2}\cos^{2}\theta m^{\prime}\right\}\,$, (B.5)
$G^{r}{}_{r}$ $=$ $-\dfrac{2r^{2}m^{\prime}}{\Sigma^{2}}\,.$ (B.6)
For the case of $\theta=\pi/2$, we find the following relations
[…]
The non-trivial components of the Einstein equations take the form
$G^{r}{}_{r}=8\pi P_{r}\,,\quad G^{\theta}{}_{\theta}=8\pi P_{\theta}\,.$ (B.11)
From these, we find
$m=M-4\pi\int drr^{2}P_{r}\,,\quad\,P_{r}=-\frac{m^{\prime}}{4\pi r^{2}}\,,\quad P_{\theta}=-\frac{m^{\prime\prime}}{8\pi r}\,.$ (B.12)
Because of the model rotation, $G^{t}_{t}$ no longer resembles a physical density. In particular, a non-vanishing cross component $T^{t}_{\varphi}=G^{t}_{\varphi}/(8\pi)$ signals the presence of an azimuthal energy flux, so that $T^{t}{}_{t}$ and $T^{\varphi}{}_{\varphi}$ are no longer the density and the pressure measured in the fluid co-moving frame. To achieve a physically meaningful value, we therefore should take the co-moving frame in which the $\{t,\varphi\}$ block of the mixed tensor is diagonal. To do this, we diagonalize the sub-matrix of $T^{\mu}{}_{\nu}$ restricted to the $\{t,\varphi\}$ subspace. Its eigenvalues are invariant and yield $T^{\prime\mu}_{\nu}=\mathrm{diag}(-\rho,\,P_{\varphi})$, where the identification relies on its associated eigenvector. The eigenvector associated to $\rho$ must be timelike, which defines the co-moving frame and the one associated to $P_{\varphi}$, spacelike, which defines the principal azimuthal direction in that frame. The values of $\rho$ and $P_{\varphi}$ in the co-moving frame therefore become
$\rho=\frac{m^{\prime}}{4\pi r^{2}}\,$, (B.13)
$P_{\varphi}=-\frac{m^{\prime\prime}}{8\pi r}\,.$ (B.14)
As the observer in the co-moving frame can be defined as $u^{\nu}=(1,0,0,0)$, therefore $\rho=\frac{1}{8\pi}G_{\mu\nu}u^{\mu}u^{\nu}$.
Appendix C Generalized Tolman–Oppenheimer–Volkoff Equation from Einstein’s Field Equations for Anisotropic Fluids
Let us consider a general spherically symmetric static metric as given in (A.1), along with a stress-energy tensor $T^{\mu}_{\nu}$ of the form (2.13).
By solviing the Einstein’s equations for the temporal and for the radial components, we obtain the following expressions
$h(r)=\left(1-\frac{2m(r)}{r}\right)^{-1},\quad f^{\prime}(r)=\frac{f(r)}{r}\left(h(r)(8\pi r^{2}P_{r}+1)-1\right).$ (C.1)
Combining these results yields an explicit expression for $f^{\prime}(r)$. We obtain equation from the radial component of the conservation of the stress-energy tensor $\nabla_{\mu}{T^{\mu}}_{r}=0$ as
$\frac{dP_{r}}{dr}=-\frac{f^{\prime}(r)}{2f(r)}(P_{r}+\rho)+\frac{2}{r}(P_{\theta}-P_{r}).$ (C.2)
Finally, substituing (C.1) into (C.2), we obtain a expression that relates tangential pressure $P_{\theta}$ and radial pressure $P_{r}$
$P_{\theta}(r)=P_{r}(r)+\frac{r}{2}P^{\prime}_{r}(r)+\frac{4\pi r^{3}P_{r}(r)+m(r)}{2(r-2m(r))}(P_{r}(r)+\rho).$ (C.3)